OpenStudy (curry):

((x+1)|x-5|)/sqrt(x-4)=>0

OpenStudy (anonymous):

\[\frac{(x + 1) |x-5|}{\sqrt{x-4}} \ge 0\] -------------------------------|----------------------------------------------------------------------- -1 Hmm |x - 5| is always positive and so is sqrt{x - 4} for x greater than 4 Hence this function is always positive for x >4 if we chose less numbers less than 4 we will have a imaginary number in the denominator not valid over real numbers

OpenStudy (anonymous):

The whole function R to R is only valid for x > 4 and for x >4 the function is always positive

OpenStudy (curry):

ty that helped so much

OpenStudy (curry):

are u going to be on for a while

OpenStudy (anonymous):

No Problem

OpenStudy (anonymous):

Yeah ask me anything but asymptotes, they don't teach it here (In India)

OpenStudy (curry):

i can if u want

OpenStudy (curry):

um look at this link it will take me too long for me explain

OpenStudy (anonymous):

Okay! :D

OpenStudy (anonymous):

Thanks!

OpenStudy (curry):

an asymptote is a line in the graph where the equation will never cross

OpenStudy (anonymous):

Hmm Interesting

OpenStudy (anonymous):

Hmm Nice the Link is good, I gotta practice asymptotes now

OpenStudy (anonymous):

Thanks Again, I will be on for an hour or so; You can Ping Me(Kick Me from Hacking Group) If help's needed